OpenStudy (anonymous):

2x^2 -1 can also be written as...?

OpenStudy (anonymous):

I did this ... and I was able to get 2(x+1)(x-1) Not sure if I'm correct.

OpenStudy (anonymous):

You basically have to just factor that buddy up there..

OpenStudy (anonymous):

Thats the only way I see it being written.

OpenStudy (anonymous):

Wait. I think Im wrong. Double checking.

OpenStudy (ranga):

^^^ you can factor 2 out if it was 2x^2 - 2 but not when it is 2x^2 - 1

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Yeahh! I just realized that -1 !!! haha.

OpenStudy (anonymous):

So therefore, you would have to find two numbers that will give you -1 but added give you 0. Then factor. Right? @ranga

OpenStudy (ranga):

Using the identity: \[\Large a ^{2} - b^{2} = (a + b)(a - b)\] we can write \[\Large 2x ^{2} - 1 = (\sqrt{2}x + 1)(\sqrt{2}x - 1)\] But that introduces radicals even though it does factorize it.

OpenStudy (anonymous):

Okay! That was way out of my league! Thought I was on the right track.... when .... I clearly wasn't. Hahaa! We learn something new everyday don't we. Lol.

OpenStudy (anonymous):

Ranga, when your completed with this, would you be up to help me with a question of mine that I wish you could help me get the answer to? @ranga

OpenStudy (ranga):

rbeckford134: sure post it on a separate thread and message me the link.

OpenStudy (anonymous):

Sounds good bud! Ill post it now. Thanks!